# Explainable AI for computational pathology identifies model limitations and tissue biomarkers

### v1

@@ Results
High attention regions drove counterintuitive effects in many specimens, while HIPPO-search-low-effect and HIPPO-search-high-effect identified more robust and consistent drivers of risk. High attention regions in high-risk cutaneous melanoma specimens (n=67) drove lower risk in 45% (n=30) of specimens. HIPPO-search-high-effect, on the other hand, identified regions that all drove higher risk and that more greatly contributed to high-risk predictions (, , independent t-test). High attention in high-risk breast cancer specimens (n=256) drove lower risk in 40% (n=102) specimens. Again, HIPPO-search-high consistently identified regions that drove higher risk in the high-risk specimens (, , independent t-test) (Fig. 5a). High attention regions in low-risk SKCM specimens (n=67) drove higher risk in 10% (n=7). HIPPO-search-low-effect identified regions that all drove lower risk and more strongly contributed to lower risk predictions (, , independent t-test). High attention regions in low-risk BRCA specimens (n=256) drove higher risk predictions in 8% (n=20) specimens. HIPPO-search-low-effect identified patches that consistently drove lower risk predictions (, , independent t-test) (Fig. 5b). This counterintuitive effect underscores that attention scores may not directly relate to model predictions. Thus, interpretations that solely rely on these features may be misguided. HIPPO search algorithms reliably identified the regions that drove risk predictions and may have value as a tool for prognostic biomarker search.
@@
Lifileucel is a promising immunotherapy for melanoma that involves isolating TILs from a patient’s tumor, replicating the TILs, and infusing them back into the patient111https://www.fda.gov/news-events/press-announcements/fda-approves-first-cellular-therapy-treat-patients-unresectable-or-metastatic-melanoma. In a phase II clinical trial, over 30% of patients responded to the therapy [66]. Identifying the patients that might respond to this therapy has the potential to improve patient outcomes and decrease costs (a single treatment may cost over $] [67]). Therefore, we sought to explore whether we could emulate this with ABMIL and HIPPO. We conducted in silico experiments to measure the effect of autologous TILs on prognosis. We used the prognostic model for cutaneous melanoma described above, and we studied the high-risk specimens in TCGA-SKCM (n=67 WSIs, n=54 patients). Counterfactuals were designed to model the injection of autologous TILs. In each specimen, TIL-positive patches were replicated , , , and (Fig. 6a). TIL-positive patches were defined using the same heuristic as above (see Methods). The change in model predictions between original specimens and autologous counterfactuals was recorded to measure the effect of additional TILs on prognosis. Cohen’s d was also calculated to quantify effect sizes. Importantly, we do not claim to demonstrate the efficacy of autologous TIL therapy through HIPPO and TCGA-SKCM. Rather, we aim to show a proof-of-principle that HIPPO may be used for hypothesis generation.
@@
Autologous TILs significantly lowered predicted risk in a dose-dependent manner. Risk decreased by () at dose (, , paired t-test), () at dose (, , paired t-test), () at dose (, , paired t-test), and () at dose (, , paired t-test) (Fig. 6b). Increasing the number of TILs by decreased predicted risk scores by over half in of high-risk specimens. Together, we demonstrated a proof-of-principle in which we use HIPPO to identify patients who may benefit from autologous TIL therapy through improved predicted prognosis following the replication of their TILs.
@@ Discussion
HIPPO’s ability to verify that models have learned biologically relevant information, as demonstrated by our analysis of TILs, is crucial for building trust in AI-driven prognostic tools. This alignment between model behavior and established biological knowledge provides a foundation for explaining model decisions to clinicians and patients, potentially facilitating the integration of AI tools into clinical workflows. It is also possible to use HIPPO’s de novo search to identify sets of patches from which expert pathologists could interpret manually to identify new biomarkers, helping to understand disease progression.
The application of HIPPO to simulate the effects of autologous TIL therapy in melanoma patients showcases the potential for in silico experimentation in computational pathology. As ABMIL models improve, this approach could have far-reaching implications for personalized medicine, offering a computational method to predict treatment responses and guide therapy selection. However, it is important to note that these simulations are based on model predictions and would require extensive clinical validation before they can be considered for patient care.
@@
While our study demonstrates the potential of HIPPO, several limitations must be acknowledged. First, the counterfactual scenarios generated by HIPPO, while informative, may not always reflect biologically plausible tissue alterations. Future work should focus on refining these interventions to more closely mimic realistic tissue changes. Second, our analysis was limited to a specific set of foundation models and datasets. Broader evaluation across diverse pathology tasks and model architectures is needed to fully characterize the generalizability of our findings. In addition the interpretations offered by HIPPO are inherently bound by the underlying model’s capabilities and potential shortcomings in representing complex biological systems.
Looking ahead, several avenues for future research emerge from this work. The integration of HIPPO with multi-modal data, including genomic and clinical information, could provide even richer insights into model behavior and biological relevance. Additionally, exploring the use of HIPPO in guiding model refinement, such as targeted fine-tuning based on identified weaknesses, represents a promising direction for improving model robustness and clinical applicability.
In conclusion, HIPPO represents a major advance in the ability to interpret AI models in computational pathology. By providing a quantitative framework for assessing the impact of specific tissue regions on model predictions, HIPPO offers a powerful tool for uncovering model limitations, verifying biological relevance, and biomarker discovery for various clinical applications. As the field of computational pathology continues to evolve, quantitative methods like HIPPO will be crucial in ensuring that AI tools are deployed responsibly and effectively in healthcare settings.
@@ Methods
HIPPO (Histopathology Interventions of Patches for Predictive Outcomes) is an explainable AI toolkit for attention-based multiple instance learning models in computational pathology. It generates counterfactual examples by manipulating whole slide image patches to explain model behavior. In ABMIL, tissue from a WSI is divided into small tiles, which are embedded using a pre-trained model. These patch embeddings serve as input to the ABMIL model, which learns to map bags of patches to specimen-level labels. HIPPO is made possible by two key features of ABMIL: (1) models are invariant to patch order, and (2) models accommodate variable number of patches. Taking advantages of these features of ABMIL, HIPPO creates counterfactual examples by adding or removing patches, allowing for evaluation of hypothetical scenarios and quantification of the effects of tissue region on model predictions. This process can be hypothesis-driven (e.g., removing tumor regions to test their necessity in predicting breast metastases), based on model attention (e.g., removing high-attention regions), or identified by greedy search (described below). Patches can also be added to measure their impact, such as introducing tumor-infiltrating lymphocytes (TILs) to TIL-deficient specimens and measuring impact on prognostic models.
HIPPO employs a search algorithm to identify patches necessary or sufficient for predictions. This algorithm iteratively measures each patch’s effect by removal and model re-evaluation. It can determine necessary patches (those causing the largest prediction drop when removed) or sufficient patches (those causing minimal change when removed). In multi-class prediction, this method can identify patches that drive the model towards specific outcomes. In regression models, the search algorithm can identify patches that either increase the output or decrease the output. The greedy search experiments in the present report were conducted on NVIDIA A100 GPUs.
@@
The patches were then embedded using five pre-trained models (embedding dimensions in parentheses): UNI (1024) [58], REMEDIS (4096) [61], Phikon (768) [62], CTransPath (768) [63], and RetCCL (2048) [64]. When embedding a slide, patches were read directly from the WSI file. A batch of 64 patches was read from the WSI, and then that batch was processed by the model to compute embeddings. The embeddings of all patches in a WSI were concatenated into one array and saved to disk for reuse. These embeddings served as inputs to all ABMIL models in the present report.

### v2

@@ Results
High attention regions drove counterintuitive effects in many specimens, while HIPPO-search-low-effect and HIPPO-search-high-effect identified more robust and consistent drivers of risk. High attention regions in high-risk cutaneous melanoma specimens (n=67) drove lower risk in 45% (n=30) of specimens. HIPPO-search-high-effect, on the other hand, identified regions that all drove higher risk and that more greatly contributed to high-risk predictions (, , independent t-test). High attention in high-risk breast cancer specimens (n=256) drove lower risk in 40% (n=102) specimens. Again, HIPPO-search-high-effect consistently identified regions that drove higher risk in the high-risk specimens (, , independent t-test) (Fig. 4a). High attention regions in low-risk SKCM specimens (n=67) drove higher risk in 10% (n=7). HIPPO-search-low-effect identified regions that all drove lower risk and more strongly contributed to lower risk predictions (, , independent t-test). High attention regions in low-risk BRCA specimens (n=256) drove higher risk predictions in 8% (n=20) specimens. HIPPO-search-low-effect identified patches that consistently drove lower risk predictions (, , independent t-test) (Fig. 4b). This counterintuitive effect underscores that attention scores may not directly relate to model predictions. Thus, interpretations that solely rely on these features may be misguided. HIPPO search algorithms reliably identified the regions that drove risk predictions and may have value as a tool for prognostic biomarker search.
@@
Lifileucel is a promising immunotherapy for melanoma that involves isolating TILs from a patient’s tumor, replicating the TILs, and infusing them back into the patient111https://www.fda.gov/news-events/press-announcements/fda-approves-first-cellular-therapy-treat-patients-unresectable-or-metastatic-melanoma. In a phase II clinical trial, over 30% of patients responded to the therapy [89]. Identifying the patients that might respond to this therapy has the potential to improve patient outcomes and decrease costs (a single treatment may cost over $] [90]). Therefore, we sought to explore whether we could emulate this with ABMIL and HIPPO. We conducted virtual experiments to measure the effect of autologous TILs on prognosis. We used the prognostic model for cutaneous melanoma described above, and we studied the high-risk specimens in TCGA-SKCM (n=67 WSIs, n=54 patients). Counterfactuals were designed to model the injection of autologous TILs. In each specimen, TIL-positive patches were replicated , , , and (Supplementary Fig. 11a). TIL-positive patches were defined using the same heuristic as above (see Methods). The change in model predictions between original specimens and autologous counterfactuals was recorded to measure the effect of additional TILs on prognosis. Cohen’s d was also calculated to quantify effect sizes. Importantly, we do not claim to demonstrate the efficacy of autologous TIL therapy through HIPPO and TCGA-SKCM. Rather, we aim to show a proof-of-principle that HIPPO may be used for hypothesis generation.
@@
Autologous TILs significantly lowered predicted risk in a dose-dependent manner. Risk decreased by () at dose (, , paired t-test), () at dose (, , paired t-test), () at dose (, , paired t-test), and () at dose (, , paired t-test) (Supplementary Fig. 11b). Increasing the number of TILs by decreased predicted risk scores by over half in of high-risk specimens. Together, we demonstrated a proof-of-principle in which we use HIPPO to identify patients who may benefit from autologous TIL therapy through improved predicted prognosis following the replication of their TILs.
HIPPO addresses limitations of attention in explaining mutation detection
To evaluate HIPPO’s advantages over attention in diagnosing misclassifications, we applied it to IDH mutation classification, a prognostic marker in gliomas [91]. Using ABMIL models trained on the EBRAINS dataset [92] (n=873), we classified wildtype (n=540) and mutant (n=333) IDH in H&E-stained specimens, employing 10-fold Monte Carlo cross-validation to assess variability. The classifiers demonstrated strong performance, with balanced accuracy of (mean std. dev.) (Fig. 5a), though the false negative (FN) rate ranged from 0.03 to 0.26 across folds. We used HIPPO and attention to identify and remove influential patches in FN specimens, aiming to increase the model’s IDH mutation probability.
When removing the top of patches identified by the strategy HIPPO-search-high-effect, predicted IDH mutation probabilities increased significantly from baseline (, independent t-test, Cohen’s ), resulting in a true positive classification for of cases. In comparison, removing the top of high-attention patches only converted FNs to true positives in of cases, with a nonsignificant effect on mutation probability (, independent t-test, Cohen’s ) (Fig. 5b). These results suggest that HIPPO was more effective than attention at identifying critical regions causing FN predictions.
While there was a large degree of overlap between the patches identified by HIPPO and attention (median: , range: to ), the patches exclusive to HIPPO were often key to resolving FN classifications. A representative sample of these patches is shown in Fig. 5c. Highlighting this, in specimen a1982c61-357f-11eb-b540-001a7dda7111, occluding high-HIPPO patches raised the IDH mutation probability from 0.24 to 0.93 (Fig. 5d), while occluding high-attention patches raised it only to 0.30 (Fig. 5e). Although HIPPO and attention overlapped by , the unique patches identified by HIPPO proved crucial in converting this FN prediction to a true positive prediction (Fig. 5e).
HIPPO also demonstrated clear advantages over attention in identifying regions that drive positive predictions. Using HIPPO-search-high-effect, we identified the patches most strongly contributing to positive predictions and compared these to patches prioritized by attention. HIPPO consistently identified more impactful patches, as shown by the steeper and more substantial decreases in model predictions when HIPPO-identified patches were removed compared to those identified by attention (Supplementary Figs. 12a and 12c).
Heatmaps generated from HIPPO outputs revealed a sharper delineation of critical regions than attention maps, which may enhance interpretability. Interestingly, HIPPO sometimes highlighted different regions than those with high attention scores, suggesting that diagnostically relevant areas may not always correspond to high-attention regions (Supplementary Figs. 12b and 12d). To further assess patch quality, we measured how many patches needed to be removed to lower a positive prediction below a probability threshold of 0.4. HIPPO required significantly fewer patches than attention (, independent t-test), indicating that it identified higher-quality, diagnostically relevant patches.
@@ Discussion
HIPPO’s ability to verify that models have learned biologically relevant information, as demonstrated by our analysis of TILs in prognostic models, is crucial for building trust in AI-driven clinical tools. This alignment between model behavior and established biological knowledge provides a foundation for explaining model decisions to clinicians and patients, potentially facilitating the integration of AI tools into clinical workflows. It is also possible to use HIPPO’s de novo search to identify sets of patches from which expert pathologists could interpret manually to identify new tissue biomarkers.
The application of HIPPO to simulate the effects of autologous TIL therapy in melanoma patients showcases the potential for virtual experimentation in computational pathology. As foundation models and ABMIL methods improve, this approach could have far-reaching implications for personalized medicine, offering a computational method to predict treatment responses and guide therapy selection. However, it is important to note that these simulations are based on model predictions and would require extensive clinical validation before they can be considered for patient care.
@@
While our study demonstrates the potential of HIPPO, several limitations must be acknowledged. First, the counterfactual scenarios generated by HIPPO, while informative, may not always reflect biologically plausible tissue alterations. Future work should focus on refining these interventions to more closely mimic realistic tissue changes. Second, HIPPO operates at the resolution of patches, which represents a fundamental limit in its ability to capture finer-grained tissue-level details. Additionally, our analysis was limited to a specific set of foundation models and datasets. Broader evaluation across diverse pathology tasks and model architectures is needed to fully characterize the generalizability of our findings. In addition the interpretations offered by HIPPO are inherently bound by the underlying model’s capabilities and potential shortcomings in representing complex biological systems.
Looking ahead, this work opens several promising directions for future research. Integrating HIPPO with multimodal data, such as genomic and clinical information, could deepen its capacity to reveal insights into both model behavior and biological relevance. While HIPPO identifies important patches de novo, interpreting these patches often requires the expertise of trained pathologists to discern the shared features driving model predictions. Alternatively, observational statistics derived from the patches, such as nuclear pleomorphism from nuclear segmentation or tissue type classifications, could help identify key features within HIPPO-identified patches. Incorporating matched spatial transcriptomics data offers another compelling avenue [93], potentially enabling biomarker discovery by linking sub-visual molecular features with morphological characteristics. Such integrations could help overcome the patch-level resolution limit of HIPPO, offering more granular insights into the interplay between molecular features and patient prognosis.
Another critical avenue involves leveraging HIPPO as a powerful auditing tool for ABMIL models. By systematically identifying model limitations, HIPPO can uncover hidden biases, weaknesses in feature representation, or instances where predictions rely on irrelevant or spurious features. This level of scrutiny can guide targeted refinements, such as fine-tuning models to correct specific deficiencies or retraining with more representative data. Moreover, HIPPO’s ability to highlight important patches provides a transparent framework for validating whether the model aligns with clinical expectations and pathology expertise. These improvements have the potential to not only enhance model robustness but also foster trust and reliability in clinical applications, paving the way for safer and more interpretable deployment in real-world settings.
In conclusion, HIPPO represents a major advance in the ability to interpret AI models in computational pathology. By providing a quantitative framework for assessing the impact of specific tissue regions on model predictions, HIPPO offers a powerful tool for uncovering model limitations, verifying biological relevance, and biomarker discovery for myriad clinical applications. As the field of computational pathology continues to evolve, quantitative methods like HIPPO will be crucial in ensuring that AI tools are deployed responsibly and effectively in healthcare settings.
@@ Methods
HIPPO (Histopathology Interventions of Patches for Predictive Outcomes) is an explainable AI toolkit for attention-based multiple instance learning (ABMIL) models in computational pathology. It generates counterfactual examples by manipulating whole slide image patches to explain model behavior. In ABMIL, tissue from a whole slide image (WSI) is divided into small tiles, which are embedded using a pre-trained model. These patch embeddings serve as input to the ABMIL model, which learns to map sets of patch embeddings to specimen-level labels. HIPPO is made possible by two key features of ABMIL: (1) models are invariant to patch order, and (2) models accommodate variable number of patches. Taking advantages of these features of ABMIL, HIPPO creates counterfactual examples by adding or removing patches, allowing for evaluation of hypothetical scenarios and quantification of the effects of tissue region on model predictions.
When annotations are available, they can guide hypothesis-driven experiments. Patches corresponding to specific annotated regions can be removed or added to test their necessity or sufficiency for model predictions. For instance, in the present study, we utilized expert annotations of tumor regions to investigate the necessity and sufficiency of tumor presence in predicting metastasis. By removing tumor patches, we assessed their impact on model predictions to determine their criticality. Conversely, tumor patches were added to other specimens to evaluate their sufficiency in driving positive predictions. Automated methods may also be used to generate annotations. In the present report, for example, regions positive for tumor-infiltrating lymphocytes were defined using an off-the-shelf cell classification network, and these regions were used to test hypotheses in prognostic models.
In many cases, annotations are unavailable. In these cases, attention-based methods can be used to select patches for experimentation. Patches with the highest attention scores, such as the top of patches, can be removed to examine their influence on predictions. The same patches can also be added to other specimens to evaluate their sufficiency for inducing specific predictions. This approach compensates for a limitation of attention mechanisms, which lack explicit information about the direction of effect. By selectively removing high-attention patches, we quantify both the direction and magnitude of their influence on model outputs.
Additionally, the HIPPO search algorithm offers an alternative for identifying critical patches. This method directly evaluates the effect of each patch on the model’s predictions, accommodating potential nonlinear interactions among patches. The algorithm operates as follows:
1.
Begin with the full set of patches for a whole slide image (WSI) and record the initial model prediction.
2.
Iteratively drop one patch at a time, measure the change in the model’s prediction, and replace the patch.
3.
Identify and permanently remove the patch that results in the lowest prediction (to identify necessary patches) or the smallest change (to identify sufficient patches).
4.
Repeat the process until all patches are exhausted or a stopping criterion is met.
The HIPPO search algorithm provides a systematic framework for identifying patches that are most critical to model predictions. This can be particularly useful when hypothesis-driven experiments reveal discrepancies between expected and actual model behavior, or when the important tissue features are not known beforehand. By identifying the regions the model prioritizes, this approach enables comparisons with domain knowledge, helps uncover unexpected decision-making patterns, and may aid in biomarker discovery. A limitation of the HIPPO search strategy is potentially long runtime, depending on the number of patches in a specimen. To address this, a batched approach can be used, wherein multiple patches are removed at each iteration. By adjusting the batch size, a trade-off can be achieved between computational efficiency and the granularity of the results.
@@
The setup for IDH mutation classification was similar to that of metastasis classification. However, CONCH [21] embeddings were used, and 10-fold Monte Carlo cross validation was employed to assess variability across models. All models were implemented in PyTorch and trained on NVIDIA RTX 2080 Ti GPUs.
@@
IDH mutation dataset
The EBRAINS dataset [92], comprising 795 patients and 873 specimens with known IDH mutation status, was used to develop IDH mutation classification models. The dataset included 540 IDH wildtype and 333 IDH mutant specimens, spanning 508 glioblastomas, 189 astrocytomas, and 176 oligodendrogliomas. All WSIs were in NDPI format with physical spacing of . Specimens were partitioned for 10-fold Monte Carlo cross-validation using the CLAM toolkit [62]. The WSIs had patches (mean standard deviation), where each patch was .
@@
The patches from CAMELYON16 were then embedded using five pre-trained models (embedding dimensions in parentheses): UNI (1024) [20], REMEDIS (4096) [16], Phikon (768) [18], CTransPath (768) [17], and RetCCL (2048) [19]. The TCGA-BRCA and TCGA-SKCM specimens for prognostic models were embedded using the UNI model, and the glioma specimens from EBRAINS were embedded using the CONCH model. When embedding a slide, patches were read directly from the WSI file. A batch of 64 patches was read from the WSI, and then that batch was processed by the model to compute embeddings. The embeddings of all patches in a WSI were concatenated into one array and saved to disk for reuse. These embeddings served as inputs to all ABMIL models in the present report.
@@
Explaining misclassifications in IDH mutation classification
We analyzed histopathology specimens with known IDH mutations but misclassified as negative by the models (i.e., false negative, FN). The objective was to evaluate whether removing the top of patches, identified by attention or HIPPO search, could increase the likelihood of a correct classification. To reduce running time of experiments, we analyzed FN specimens with fewer than patches (n=38). For each specimen, the softmax probability for IDH mutation was recorded. Next, the of patches with the highest attention scores were removed, and the updated model probability was documented. In addition, the top of patches identified by HIPPO-search-high-effect were removed, and the updated model probability was recorded (search algorithm described below). The final IDH mutation probabilities after the different methods of patch removal were compared across baseline, attention, and HIPPO conditions. Cohen’s d was used to assess effect sizes, and independent t-tests evaluated statistical significance between conditions. The experiment was conducted exclusively on specimens from the test set of each fold during the 10-fold Monte Carlo cross-validation, ensuring that no specimens from the training or validation sets were included.
The HIPPO-search-high-effect algorithm iteratively removed individual patches, replacing each patch after recording its impact on mutation probability. The patch whose removal most increased the probability of IDH mutation was permanently excluded, with this process repeated until of patches were removed. The search was performed using NVIDIA A100 GPUs.
Identifying regions driving positive IDH mutation prediction
We evaluated the extent to which attention and HIPPO search identified patches driving positive predictions, focusing on specimens with positive predictions in the first Monte Carlo cross-validation fold (n=25). The HIPPO-search-high-effect algorithm was applied, removing patches in batches of 10 per iteration to accelerate runtime. Patches were ranked by their impact on IDH mutation probabilities, with the 10 patches whose removal most reduced the prediction probability excluded at each step. The IDH mutation probability was recorded after each batch, continuing until of patches were removed. A similar approach was used for attention, removing patches in descending order of attention scores in batches of 10 and recording model outputs after each removal. Model outputs for each specimen were visualized as line plots, comparing results from HIPPO and attention. Heatmaps of attention and HIPPO outputs were viewed using QuPath for qualitative assessment.
To assess the efficiency of HIPPO in identifying critical patches, we conducted an experiment to determine how many patches needed to be removed to reduce the model probability below 0.4, hypothesizing that HIPPO would require fewer patches than attention due to its ability to better capture the most important regions driving predictions. The HIPPO and attention outputs from above were used in this experiment. The ratio of patches required (i.e., number of patches required normalized by number of patches in specimen) were compared between attention and HIPPO using independent t-test and visualized using a scatter plot.